OpenStudy (anonymous):

I have $6250 and open a checking and a CD. The bank pays 2.6% interest on CD and 0.05% ineterest on checking. At end of year I earn $127.50 how much was made on each account

OpenStudy (amistre64):

c = 6250-s c(.026) + s(.0005) = 127.50 (6250-s)(.026) + s(.0005) = 127.50 162.50 - .026s + .0005s = 127.50 162.50 - 127.50 = .026s -.0005s 35 = 0.0255s 35 / .0255 = s initial investment in savings = 1372.55 initial investment on CD = 4877.45 CD = 4877.45(.026) = 126.8137 savings = 1372.55(.0005) = 0.686275 together they equal: 127.499975

OpenStudy (amistre64):

almost lost all that valueable work to a frozen screen.... but I is smart :)